### v1

@@ 4.3 Non-informative cluster size and a binary outcome
Table 3: Simulation results under a data generating process with non-informative cluster size and a binary outcome. GLM: logistic generalized linear model based on cluster-level means; GLMM: logistic generalized linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Monte Carlo standard errors for performance metric estimates are presented in Web Table 3.
$m$ Covariate Outcome model Method BIAS MCSD AESE COV BIAS MCSD AESE COV
adjustment
@@
[…]
MRS 0.9 0.32 0.32 95.1 2.3 0.35 0.35 95.0
[…]
MRS 0.9 0.32 0.32 95.1 2.3 0.35 0.35 94.7
[…]
MRS 0.9 0.32 0.32 95.2 2.3 0.35 0.35 94.9
[…]
MRS 0.7 0.32 0.33 95.4 2.1 0.35 0.36 94.2
[…]
MRS 0.7 0.19 0.24 96.2 1.0 0.19 0.24 96.5
[…]
MRS 0.5 0.18 0.21 95.3 0.8 0.18 0.21 95.9
[…]
MRS 0.4 0.18 0.22 95.2 0.8 0.18 0.22 95.9
[…]
MRS 0.5 0.18 0.21 95.5 0.9 0.18 0.22 95.6
@@
[…]
MRS -6.0 0.19 0.18 92.3 -3.5 0.20 0.19 92.7
[…]
MRS -5.8 0.19 0.18 92.1 -3.4 0.20 0.19 92.7
[…]
MRS -5.9 0.19 0.18 92.3 -3.5 0.20 0.19 92.8
[…]
MRS -5.8 0.19 0.18 92.1 -3.4 0.20 0.19 92.7
[…]
MRS -3.3 0.12 0.12 96.1 -1.9 0.11 0.12 96.4
[…]
MRS -3.1 0.11 0.12 95.0 -1.8 0.11 0.12 96.0
[…]
MRS -3.2 0.11 0.12 95.3 -1.9 0.11 0.12 95.8
[…]
MRS -3.2 0.12 0.12 95.3 -1.9 0.11 0.12 95.7
@@ 4.4 Informative cluster size and a binary outcome
Table 4: Simulation results under a data generating process with informative cluster size and a binary outcome. GLM: logistic generalized linear model based on cluster-level means; GLMM: logistic generalized linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Monte Carlo standard errors for performance metric estimates are presented in Web Table 4.
$m$ Covariate Outcome model Method BIAS MCSD AESE COV BIAS MCSD AESE COV
adjustment
@@
$30$$\times$ (W5) GLM Coef 1.3 0.36 0.36 95.5 -25.9 0.36 0.36 85.2
MRS 1.3 0.36 0.36 95.6 0.3 0.41 0.42 95.4
(W6) GLMM Coef 31.6 0.45 0.43 90.0 -3.7 0.45 0.43 93.1
MRS 1.3 0.36 0.36 95.8 0.1 0.41 0.42 95.3
(W7) GEE-exch Coef 3.2 0.36 0.37 95.5 -24.5 0.36 0.37 86.9
MRS 1.3 0.36 0.36 95.7 0.3 0.41 0.42 95.4
(W8) GEE-ind Coef 37.0 0.40 0.42 89.8 0.3 0.40 0.42 95.6
MRS 1.3 0.37 0.37 95.4 0.3 0.40 0.42 95.5
$\checkmark$ (W5) GLM Coef 25.9 0.32 0.29 86.0 -7.9 0.32 0.29 89.7
MRS 3.6 0.25 0.34 97.0 -4.1 0.26 0.36 97.8
(W6) GLMM Coef 39.6 0.32 0.27 72.0 2.1 0.32 0.27 89.2
MRS 3.6 0.25 0.30 96.3 -4.3 0.26 0.31 96.6
(W7) GEE-exch Coef 27.9 0.30 0.36 90.7 -6.4 0.30 0.36 96.0
MRS 3.3 0.25 0.29 96.1 -4.4 0.26 0.31 96.7
(W8) GEE-ind Coef 58.4 0.31 0.39 75.2 15.9 0.31 0.39 94.2
MRS 4.3 0.25 0.30 96.3 -3.2 0.26 0.32 96.2
@@
$100$$\times$ (W5) GLM Coef -0.8 0.19 0.19 96.4 -27.4 0.19 0.19 72.8
MRS -0.8 0.19 0.19 96.0 -0.9 0.22 0.22 95.9
(W6) GLMM Coef 28.3 0.23 0.23 85.5 -6.1 0.23 0.23 94.5
MRS -0.8 0.19 0.19 96.0 -0.9 0.22 0.22 95.8
(W7) GEE-exch Coef 5.1 0.19 0.19 95.7 -23.1 0.19 0.19 79.6
MRS -0.9 0.19 0.19 96.0 -0.9 0.22 0.22 96.0
(W8) GEE-ind Coef 35.1 0.21 0.22 77.7 -1.1 0.21 0.22 96.0
MRS -1.1 0.19 0.19 96.3 -1.1 0.21 0.21 96.0
$\checkmark$ (W5) GLM Coef 20.8 0.16 0.15 81.2 -11.6 0.16 0.15 85.2
MRS 1.5 0.13 0.14 96.6 -1.1 0.13 0.15 95.4
(W6) GLMM Coef 41.4 0.16 0.15 49.2 3.5 0.16 0.15 93.4
MRS 1.4 0.13 0.14 96.2 -1.2 0.13 0.14 94.9
(W7) GEE-exch Coef 32.8 0.15 0.16 69.7 -2.8 0.15 0.16 95.6
MRS 1.4 0.13 0.14 96.0 -1.1 0.13 0.14 94.8
(W8) GEE-ind Coef 57.7 0.15 0.16 26.2 15.4 0.15 0.16 86.2
MRS 1.7 0.13 0.14 96.1 -0.8 0.13 0.14 94.7
@@ 4.5 Additional simulations with a smaller sample size and under constrained randomization
We also replicate the above simulation design under informative cluster size (I) when the number of clusters is small and (II) when the study adopts covariate-constrained randomization. The additional simulations under (I) are used to explore whether MRS remains robust to estimate cluster-ATE and individual-ATE when $m$ is as small as $12$. The additional simulations under (II) are used to explore whether MRS retains its favorable performance characteristics over Coef under a more complex randomization scheme. Detailed configurations are provided in Web Appendix D, with full numerical results summarized in Web Tables 10-15. Overall, when $m\leq 20$, MRS remains robust to the choice of outcome model, but the jackknife approach can sometimes yield conservative inference. Next, under covariate-constrained randomization, the MRS estimators continue to exhibit negligible bias and maintain nominal or slightly conservative coverage. In contrast, the performance of the Coef estimators remains sensitive to the choice of working outcome model, and constrained randomization does not mitigate the potential bias inherent to certain Coef estimators under informative cluster size.
@@ 4.6 Testing for informative cluster size
[…]
@@ 5 An illustrative data example
The data analysis results for the PPACT trial are summarized in Figure 1. Both cluster-level ($\Delta_{C}$) and individual-level ($\Delta_{I}$) estimators are calculated using the model-robust standardization estimator, with and without covariate adjustment. The model-based coefficient estimates are reproduced under both $\widehat{\Delta}_{C}$ and $\widehat{\Delta}_{I}$. The results show a clear pattern. Regardless of covariate adjustment, the treatment coefficient from different regression models can be somewhat different, whereas the model-robust standardization estimators tend to be much less sensitive to the choice of working outcome model. Under informative cluster size, previous studies13, 11, 5, 4 and our simulations in Section 4.2 have shown that, regardless of covariate adjustment, treatment effect coefficients from (W1) and (W4) are consistent to $\widehat{\Delta}_{C}$ and $\widehat{\Delta}_{I}$, respectively. This is also indirectly supported by our data analysis, as the treatment coefficient point estimates in those cases are identical to those under model-robust standardization. Furthermore, under model-robust standardization, we observe that covariate adjustment leads to some differences in the point estimates, likely due to addressing residual imbalance in finite samples. For each working regression model, covariate adjustment consistently results in narrower confidence intervals and leads to efficiency gain compared to unadjusted analyses. To conclude, based on model-robust standardization, the intervention leads to notable reduction of the PEGS score at one year, based on either the cluster-ATE ($\Delta_{C}$) or the individual-ATE ($\Delta_{I}$) estimates.
@@ 6 Discussion
Although there have been three decades of methods development for analyzing CRTs, a stronger call for clarifying treatment effect estimands has only emerged in the recent literature.4, 5 In that context, several prior studies have examined whether the model-based treatment effect coefficient can target marginal estimands under the potential outcomes framework. For example, Wang et al.22 proved that without informative cluster size, the treatment coefficient estimator from a linear mixed model is consistent to the average treatment effect estimand (also see Table 1 for an empirical verification) under arbitrary working model misspecification. However, that estimator generally converges to a less interpretable target parameter under informative cluster sizes.11, 5 On the other hand, as two examples of model-assisted procedures, the treatment coefficient estimator from cluster-level linear regression targets the cluster-ATE, and that from the (unweighted) independence GEE targets individual-ATE, under arbitrary model misspecification.13, 4 Although not examined by our simulations, cluster-level linear regression can also target individual-ATE and independence GEE can target individual-ATE once appropriate weights are specified.13, 11, 4, 5 While these insightful results provide timely and important clarifications, a caveat is that they have only been developed for specific regression models. In addition, although the cluster-level linear regression and independence GEE are robust to arbitrary misspecification, they are much less commonly used in practice compared to mixed model regression for analyzing CRTs.19Given this discrepancy, an important objective of our paper is to demonstrate that there exists a unifying procedure—model-robust standardization—that can incorporate any regression models (including generalized linear mixed models and exchangeable GEE) previously developed for CRTs and still consistently target cluster-ATE or individual-ATE as two typical marginal estimands. This has important practical implications because one can automatically achieve estimand-aligned inference in CRTs using their preferred model specification as a working model in our procedure. To facilitate the application of this approach in practice, we also provide an R package MRStdCRT that is available at https://github.com/deckardt98/MRStdCRT. A short tutorial is available in Appendix B to explain the syntax and output with the example dataset analyzed in Section 5.
Given a common set of pre-specified baseline covariates, although the choices of outcome models do not affect consistency for a model-robust standardization approach, they can affect statistical efficiency. In fact, the model-robust standardization approach provides a transparent framework for comparing efficiency across different outcome modeling methods (e.g., generalized linear mixed models and GEE) by targeting a common estimand. Future simulation studies can thus be conducted within this framework to identify best-performing outcome models under different data generating processes, particularly for ratio estimands and binary outcomes. Furthermore, integrating data-adaptive machine learning methods to estimate ${E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ in this framework can theoretically improve asymptotic efficiency compared to misspecified parametric models. Since the primary aim of this paper is to introduce the model-robust standardization approach to improve parametric or semi-parametric regression modeling practices, we refer to Wang et al.18, 17, 52, Balzer et al.15, Benitez et al.10, and Nugent et al.16 for theoretical and practical considerations on applying machine learning estimators in CRTs. Lastly, although machine learning methods are appealing with large numbers of clusters, further clarifying their relative strengths and limitations with smaller cluster sizes remains valuable for guiding practice.
@@
While the cluster jackknife variance estimator offers a computationally convenient recipe for inference under model-robust standardization, for any given outcome regression model, our simulation findings highlight important opportunities for future research to improve this estimator. In Section 4, even though the jackknife variance estimator is generally close to the Monte Carlo variance and leads to nominal coverage with $m=100$ clusters, it becomes slightly conservative as the number of clusters decreases to $m=30$, and may be increasingly conservative with even fewer clusters (Web Appendix D). This pattern aligns with earlier findings in the context of simple linear regression with clustered data, as documented by Hansen.46 To some extent, our empirical results extend their observations to a broader class of outcome regression models. Although small-sample corrections for model-based inference in CRTs have been more extensively studied,47, 50, 51, 48 analogous adjustments for model-robust standardization are yet to be more systematically investigated. There remains substantial uncertainty about how best to improve confidence interval coverage for estimating cluster-ATE and individual-ATE in a small CRT. Therefore, developing principled small-sample corrections that improve the cluster jackknife represents an important and largely open area for future research. Synergistically, small-sample adjustments may also help mitigate the conservativeness observed for tests for informative cluster size.
@@
This work stimulates the following areas that we plan to pursue in future research. First, Bayesian methods have been developed for the analysis of CRTs, but model-robust inference with Bayesian hierarchical models has not been fully elucidated. It will be interesting to develop the model-robust analogue when ${E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ is informed by posterior distributions of model parameters. Second, we have not study survival outcomes, but plan to further generalize the model-robust standardization estimator to accommodate covariate-dependent right censoring. Third, an increasing number of CRTs are designed with outcome data collected over several different time periods, and additionally with treatment rolled out in a staggered fashion. Specifically in stepped wedge CRTs, treatment effect estimands under the potential outcomes framework have been formalized recently, when the cluster-period size is informative53 and when the treatment effect can depend on treatment duration.54 It would be useful to generalize the model-robust standardization estimator such that estimand-aligned analysis of SW-CRTs can be achieved with an arbitrary outcome regression model. Fourth, while our focus is on marginal estimands defined under the potential outcomes framework, we acknowledge that these may not be the only estimands of interest. In particular, cluster-specific and individual-specific estimands, though not addressed in this work, can also be important for informing decisions at the level of specific cluster stakeholders or individual patients and are useful to study in future work. Finally, we have not explored the implications of missing outcomes within the model-robust standardization framework. A practical approach is to combine model-robust standardization with multilevel multiple imputation, consistent with existing recommendations for handling missing outcomes in model-based analyses of CRTs.55 However, a comprehensive investigation of this strategy is left for future research.

### v2

@@ 4.3 Non-informative cluster size and a binary outcome
Table 3: Simulation results under a data generating process with non-informative cluster size and a binary outcome. GLM: logistic generalized linear model based on cluster-level means; GLMM: logistic generalized linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Each empirical coverage probability is reported together with its Monte Carlo 95% confidence limits (95% CL). Additional Monte Carlo standard errors for performance metric estimates are presented in Web Table 3.
$m$ Covariate Outcome model Method BIAS MCSD COV BIAS MCSD COV
adjustment (AESE) (95% CL) (AESE) (95% CL)
@@
$30$ $\times$ (W1) LM Coef 0.9 0.32 (0.32) 95.5 (94.5, 96.5) \ \ \
MRS 0.9 0.32 (0.32) 95.1 (94.1, 96.1) 2.3 0.35 (0.35) 95.0 (94.0, 96.0)
(W2) LMM Coef 19.2 0.37 (0.35) 92.8 (91.5, 94.1) \ \ \
MRS 0.9 0.32 (0.32) 95.1 (94.1, 96.1) 2.3 0.35 (0.35) 94.7 (93.7, 95.7)
(W3) GEE-exch Coef 1.1 0.32 (0.33) 95.7 (94.7, 96.7) \ \ \
MRS 0.9 0.32 (0.32) 95.2 (94.2, 96.2) 2.3 0.35 (0.35) 94.9 (93.9, 95.9)
(W4) GEE-ind Coef 2.1 0.35 (0.36) 94.5 (93.3, 95.7) \ \ \
MRS 0.7 0.32 (0.33) 95.4 (94.4, 96.4) 2.1 0.35 (0.36) 94.2 (93.1, 95.3)
$\checkmark\checkmark$ (W1) LM Coef 16.8 0.22 (0.20) 89.7 (88.1, 91.3) \ \ \
MRS 0.7 0.19 (0.24) 96.2 (95.2, 97.2) 1.0 0.19 (0.24) 96.5 (95.6, 97.4)
(W2) LMM Coef 19.3 0.21 (0.18) 84.9 (83.1, 86.7) \ \ \
MRS 0.5 0.18 (0.21) 95.3 (94.3, 96.3) 0.8 0.18 (0.21) 95.9 (95.0, 96.8)
(W3) GEE-exch Coef 16.2 0.21 (0.24) 93.6 (92.4, 94.8) \ \ \
MRS 0.5 0.18 (0.22) 95.2 (94.2, 96.2) 0.8 0.18 (0.22) 95.9 (95.0, 96.8)
(W4) GEE-ind Coef 16.5 0.21 (0.25) 94.0 (92.9, 95.1) \ \ \
MRS 0.5 0.18 (0.21) 95.5 (94.6, 96.4) 0.9 0.18 (0.22) 95.6 (94.6, 96.6)
@@
$100$ $\times$ (W1) LM Coef -6.0 0.19 (0.18) 92.2 (91.0, 93.4) \ \ \
MRS -6.0 0.19 (0.18) 92.3 (91.1, 93.5) -3.5 0.20 (0.19) 92.7 (91.5, 93.9)
(W2) LMM Coef 10.2 0.22 (0.20) 92.4 (91.2, 93.6) \ \ \
MRS -5.8 0.19 (0.18) 92.1 (90.9, 93.3) -3.4 0.20 (0.19) 92.7 (91.5, 93.9)
(W3) GEE-exch Coef -5.1 0.19 (0.18) 92.6 (91.4, 93.8) \ \ \
MRS -5.9 0.19 (0.18) 92.3 (91.1, 93.5) -3.5 0.20 (0.19) 92.8 (91.6, 94.0)
(W4) GEE-ind Coef -3.4 0.20 (0.19) 93.2 (92.0, 94.4) \ \ \
MRS -5.8 0.19 (0.18) 92.1 (90.9, 93.3) -3.4 0.20 (0.19) 92.7 (91.5, 93.9)
$\checkmark\checkmark$ (W1) LM Coef 11.1 0.13 (0.13) 90.7 (89.2, 92.2) \ \ \
MRS -3.3 0.12 (0.12) 96.1 (95.2, 97.0) -1.9 0.11 (0.12) 96.4 (95.5, 97.3)
(W2) LMM Coef 14.8 0.13 (0.13) 87.4 (85.7, 89.1) \ \ \
MRS -3.1 0.11 (0.12) 95.0 (94.0, 96.0) -1.8 0.11 (0.12) 96.0 (95.1, 96.9)
(W3) GEE-exch Coef 11.4 0.12 (0.13) 92.5 (91.2, 93.8) \ \ \
MRS -3.2 0.11 (0.12) 95.3 (94.4, 96.2) -1.9 0.11 (0.12) 95.8 (94.9, 96.7)
(W4) GEE-ind Coef 12.0 0.13 (0.14) 92.5 (91.2, 93.8) \ \ \
MRS -3.2 0.12 (0.12) 95.3 (94.4, 96.2) -1.9 0.11 (0.12) 95.7 (94.8, 96.6)
@@ 4.4 Informative cluster size and a binary outcome
Table 4: Simulation results under a data generating process with informative cluster size and a binary outcome. GLM: logistic generalized linear model based on cluster-level means; GLMM: logistic generalized linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Each empirical coverage probability is reported together with its Monte Carlo 95% confidence limits (95% CL). Additional Monte Carlo standard errors for performance metric estimates are presented in Web Table 4.
$m$ Covariate Outcome model Method BIAS MCSD COV BIAS MCSD COV
adjustment (AESE) (95% CL) (AESE) (95% CL)
@@
$30$ $\times$ (W1) LM Coef 1.3 0.36 (0.36) 95.5 (94.5, 96.5) -25.9 0.36 (0.36) 85.2 (83.3, 87.1)
MRS 1.3 0.36 (0.36) 95.6 (94.6, 96.6) 0.3 0.41 (0.42) 95.4 (94.4, 96.4)
(W2) LMM Coef 31.6 0.45 (0.43) 90.0 (88.4, 91.6) -3.7 0.45 (0.43) 93.1 (91.9, 94.3)
MRS 1.3 0.36 (0.37) 95.8 (94.8, 96.8) 0.1 0.41 (0.42) 95.3 (94.3, 96.3)
(W3) GEE-exch Coef 3.2 0.36 (0.37) 95.5 (94.5, 96.5) -24.5 0.36 (0.37) 86.9 (85.1, 88.7)
MRS 1.3 0.36 (0.36) 95.7 (94.7, 96.7) 0.3 0.41 (0.42) 95.4 (94.4, 96.4)
(W4) GEE-ind Coef 37.0 0.40 (0.42) 89.8 (88.2, 91.4) 0.3 0.40 (0.42) 95.6 (94.6, 96.6)
MRS 1.3 0.37 (0.37) 95.4 (94.4, 96.4) 0.3 0.40 (0.42) 95.5 (94.5, 96.5)
$\checkmark\checkmark$ (W1) LM Coef 25.9 0.32 (0.29) 86.0 (84.2, 87.8) -7.9 0.32 (0.29) 89.7 (88.1, 91.3)
MRS 3.6 0.25 (0.34) 97.0 (96.1, 97.9) -4.1 0.26 (0.36) 97.8 (97.0, 98.6)
(W2) LMM Coef 39.6 0.32 (0.27) 72.0 (69.6, 74.4) 2.1 0.32 (0.27) 89.2 (87.6, 90.8)
MRS 3.6 0.25 (0.30) 96.3 (95.4, 97.2) -4.3 0.26 (0.31) 96.6 (95.7, 97.5)
(W3) GEE-exch Coef 27.9 0.30 (0.36) 90.7 (89.2, 92.2) -6.4 0.30 (0.36) 96.0 (95.1, 96.9)
MRS 3.3 0.25 (0.29) 96.1 (95.2, 97.0) -4.4 0.26 (0.31) 96.7 (95.8, 97.6)
(W4) GEE-ind Coef 58.4 0.31 (0.39) 75.2 (72.9, 77.5) 15.9 0.31 (0.39) 94.2 (93.2, 95.2)
MRS 4.3 0.25 (0.30) 96.3 (95.4, 97.2) -3.2 0.26 (0.32) 96.2 (95.3, 97.1)
@@
$100$ $\times$ (W1) LM Coef -0.8 0.19 (0.19) 96.4 (95.5, 97.3) -27.4 0.19 (0.19) 72.8 (70.4, 75.2)
MRS -0.8 0.19 (0.19) 96.0 (95.0, 97.0) -0.9 0.22 (0.22) 95.9 (94.9, 96.9)
(W2) LMM Coef 28.3 0.23 (0.23) 85.5 (83.7, 87.3) -6.1 0.23 (0.23) 94.5 (93.4, 95.6)
MRS -0.8 0.19 (0.19) 96.0 (95.0, 97.0) -0.9 0.22 (0.22) 95.8 (94.8, 96.8)
(W3) GEE-exch Coef 5.1 0.19 (0.19) 95.7 (94.8, 96.6) -23.1 0.19 (0.19) 79.6 (77.5, 81.7)
MRS -0.9 0.19 (0.19) 96.0 (95.0, 97.0) -0.9 0.22 (0.22) 96.0 (95.1, 96.9)
(W4) GEE-ind Coef 35.1 0.21 (0.22) 77.7 (75.5, 79.9) -1.1 0.21 (0.22) 96.0 (95.1, 96.9)
MRS -1.1 0.19 (0.19) 96.3 (95.4, 97.2) -1.1 0.21 (0.21) 96.0 (95.1, 96.9)
$\checkmark\checkmark$ (W1) LM Coef 20.8 0.16 (0.15) 81.2 (79.1, 83.3) -11.6 0.16 (0.15) 85.2 (83.3, 87.1)
MRS 1.5 0.13 (0.14) 96.6 (95.8, 97.4) -1.1 0.13 (0.15) 95.4 (94.5, 96.3)
(W2) LMM Coef 41.4 0.16 (0.15) 49.2 (46.5, 51.9) 3.5 0.16 (0.15) 93.4 (92.2, 94.6)
MRS 1.4 0.13 (0.14) 96.2 (95.4, 97.0) -1.2 0.13 (0.14) 94.9 (94.0, 95.8)
(W3) GEE-exch Coef 32.8 0.15 (0.16) 69.7 (67.4, 72.0) -2.8 0.15 (0.16) 95.6 (94.8, 96.4)
MRS 1.4 0.13 (0.14) 96.0 (95.2, 96.8) -1.1 0.13 (0.14) 94.8 (93.9, 95.7)
(W4) GEE-ind Coef 57.6 0.15 (0.16) 26.2 (23.8, 28.6) 15.4 0.15 (0.16) 86.2 (84.3, 88.1)
MRS 1.7 0.13 (0.14) 96.1 (95.3, 96.9) -0.8 0.13 (0.14) 94.7 (93.8, 95.6)
@@ 4.5 Additional simulations with a smaller sample size and under constrained randomization
We also replicate the above simulation design under informative cluster size (I) when the number of clusters is small and (II) when the study adopts covariate-constrained randomization. The additional simulations under (I) are used to explore whether MRS remains robust to estimate cluster-ATE and individual-ATE when $m$ is as small as $12$. The additional simulations under (II) are used to explore whether MRS retains its favorable performance characteristics over Coef under a more complex randomization scheme. Detailed configurations are provided in Web Appendix E, with full numerical results summarized in Web Tables 6-11. Overall, when $m\leq 20$, MRS remains robust to the choice of outcome model, but the jackknife approach can sometimes yield conservative inference. Next, under covariate-constrained randomization, the MRS estimators continue to exhibit negligible bias and maintain nominal or slightly conservative coverage. In contrast, the performance of the Coef estimators remains sensitive to the choice of working outcome model, and constrained randomization does not mitigate the potential bias inherent to certain Coef estimators under informative cluster size.
@@ 4.6 Testing for informative cluster size
$Y_{ij}(a)\sim\mathcal{N}\left(\frac{H_{1i}X_{1ij}^{2}}{5{E}[N_{i}]}+\cos(H_{2i})X_{2ij}+|H_{2i}|\sin(X_{2ij})+\left\{\frac{\delta N_{i}^{2}\log(N_{i})}{({E}[N_{i}])^{2}}+1\right\}a+\gamma_{i}a,\,1\right),~~~~a\in\{0,1\}.$
@@ 5 An illustrative data example
The data analysis results for the PPACT trial are summarized in Figure 1. Both cluster-level ($\Delta_{C}$) and individual-level ($\Delta_{I}$) estimators are calculated using the model-robust standardization estimator, with and without covariate adjustment. The model-based coefficient estimates are reproduced under both $\widehat{\Delta}_{C}$ and $\widehat{\Delta}_{I}$. The results show a clear pattern. Regardless of covariate adjustment, the treatment coefficient from different regression models can be somewhat different, whereas the model-robust standardization estimators tend to be much less sensitive to the choice of working outcome model. Under informative cluster size, previous studies12, 11, 5, 4 and our simulations in Section 4.2 have shown that, regardless of covariate adjustment, treatment effect coefficients from (W1) and (W4) are consistent to $\widehat{\Delta}_{C}$ and $\widehat{\Delta}_{I}$, respectively. This is also indirectly supported by our data analysis, as the treatment coefficient point estimates in those cases are identical to those under model-robust standardization. Furthermore, under model-robust standardization, we observe that covariate adjustment leads to some differences in the point estimates, likely due to addressing residual imbalance in finite samples. For each working regression model, covariate adjustment consistently results in narrower confidence intervals and leads to efficiency gain compared to unadjusted analyses. To conclude, based on model-robust standardization, the intervention leads to notable reduction of the PEGS score at one year, based on either the cluster-ATE ($\Delta_{C}$) or the individual-ATE ($\Delta_{I}$) estimates.
@@ 6 Discussion
Although there have been three decades of methods development for analyzing CRTs, a stronger call for clarifying treatment effect estimands has only emerged in the recent literature.4, 5 In that context, several prior studies have examined whether the model-based treatment effect coefficient can target marginal estimands under the potential outcomes framework. For example, Wang et al.21 proved that without informative cluster size, the treatment coefficient estimator from a linear mixed model is consistent to the average treatment effect estimand (also see Table 1 for an empirical verification) under arbitrary working model misspecification. However, that estimator generally converges to a less interpretable target parameter under informative cluster sizes.11, 5 On the other hand, as two examples of model-assisted procedures, the treatment coefficient estimator from cluster-level linear regression targets the cluster-ATE, and that from the (unweighted) independence GEE targets individual-ATE, under arbitrary model misspecification.12, 4 Although not examined by our simulations, cluster-level linear regression can also target individual-ATE and independence GEE can target individual-ATE once appropriate weights are specified.12, 11, 4, 5 While these insightful results provide timely and important clarifications, a caveat is that they have only been developed for specific regression models. In addition, although the cluster-level linear regression and independence GEE are robust to arbitrary misspecification, they are much less commonly used in practice compared to mixed model regression for analyzing CRTs.18Given this discrepancy, an important objective of our paper is to demonstrate that there exists a unifying procedure—model-robust standardization—that can incorporate any regression models (including generalized linear mixed models and exchangeable GEE) previously developed for CRTs and still consistently target cluster-ATE or individual-ATE as two typical marginal estimands. This has important practical implications because one can automatically achieve estimand-aligned inference in CRTs using their preferred model specification as a working model in our procedure. We acknowledge that this work primarily focuses on aligning estimation procedures with target estimands, rather than on maximizing statistical efficiency. While covariate adjustment can generally improve efficiency relative to unadjusted estimators, such gains are not specific to the proposed model-robust standardization approach. For broader discussions on leveraging flexible outcome models—including machine learning methods—for maximizing efficiency gains in CRTs, we refer readers to recent work by Wang et al.,16, 17, 50 Balzer et al.,14 Benitez et al.,10 and Nugent et al.15 Finally, to facilitate the application of our model-robust standardization approach in practice, we also provide an R package MRStdCRT that is available at https://github.com/deckardt98/MRStdCRT. A short tutorial is available in Appendix B to explain the syntax and output with the example dataset analyzed in Section 5.
@@
While the cluster jackknife variance estimator offers a computationally convenient recipe for inference under model-robust standardization, our simulation findings highlight important opportunities for future research to improve this estimator. In Section 4, even though the jackknife variance estimator is generally close to the Monte Carlo variance and leads to nominal coverage with $m=100$ clusters, it becomes slightly conservative as the number of clusters decreases to $m=30$, and may be increasingly conservative with even fewer clusters (Web Appendix E). This pattern aligns with earlier findings in the context of simple linear regression with clustered data, as documented by Hansen.45 To some extent, our empirical results extend their observations to a broader class of outcome regression models. Although small-sample corrections for model-based inference in CRTs have been more extensively studied,46, 49 analogous adjustments for model-robust standardization are yet to be more systematically investigated. There remains substantial uncertainty about how best to improve confidence interval coverage for estimating cluster-ATE and individual-ATE in a small CRT. Therefore, developing principled small-sample corrections that improve the cluster jackknife represents an important and largely open area for future research. Synergistically, small-sample adjustments may also help mitigate the conservativeness observed for tests for informative cluster size.
@@
This work stimulates the following areas that we plan to pursue in future research. First, Bayesian methods have been developed for the analysis of CRTs, but model-robust inference with Bayesian hierarchical models has not been fully elucidated. It will be interesting to develop the model-robust analogue when ${E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ is informed by posterior distributions of model parameters. Second, we have not study survival outcomes, but plan to further generalize the model-robust standardization estimator to accommodate covariate-dependent right censoring. Third, an increasing number of CRTs are designed with outcome data collected over several different time periods, and additionally with treatment rolled out in a staggered fashion. Specifically in stepped wedge CRTs, treatment effect estimands under the potential outcomes framework have been formalized recently, when the cluster-period size is informative51 and when the treatment effect can depend on treatment duration.52 It would be useful to generalize the model-robust standardization estimator such that estimand-aligned analysis of SW-CRTs can be achieved with an arbitrary outcome regression model. Fourth, while our focus is on marginal estimands defined under the potential outcomes framework, we acknowledge that these may not be the only estimands of interest. In particular, cluster-specific and individual-specific estimands, though not addressed in this work, can also be important for informing decisions at the level of specific cluster stakeholders or individual patients and are useful to study in future work. Finally, we have not explored the implications of missing outcomes within the model-robust standardization framework. A practical approach is to combine model-robust standardization with multilevel multiple imputation, consistent with existing recommendations for handling missing outcomes in model-based analyses of CRTs.53 However, a comprehensive investigation of this strategy is left for future research.